# Does it matter? A more careful treatment of density fluctuations in 21-cm simulations

### v1

@@ II.1 Matter density fields at the dark ages
During the dark ages, the fluctuations in the matter density field were still much smaller than unity and linear perturbation theory can be applied in that epoch. After recombination took place, on sub-horizon scales larger than the Jeans scale, the linearized continuity and Euler equations for the baryon and CDM fluids are identical Baumann (2022); Dodelson (2003),
@@
$\dot{\mathbf{v}}_{b}+H\mathbf{v}_{b}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ 00 0 (2)
@@
where $\delta_{b}$ ($\delta_{c}$) is the baryon (CDM) energy over-density, $\mathbf{v}_{b}$ ($\mathbf{v}_{c}$) is the proper baryon (CDM) peculiar velocity, $a$ is the scale factor, and $H\equiv\dot{a}/a$ is the Hubble parameter. Overdots denote derivatives with respect to the cosmological time $t$. Eqs. (1)-(4) are accompanied with the Poisson equation for the gravitational potential $\Phi$,
@@
We can combine Eqs. (1)-(6) into two coupled second-order differential equations of the same form ($i=b,c$ for baryons and CDM),
@@
The CDM SDGF is nearly scale-invariant and can be very well approximated by $\mathcal{D}_{c}\left(k,z\right)\approx D\left(z\right)$, apart for BAO features that appear at $z\lesssim 1$ (while adopting the SIGF for the CDM density field causes “fake BAOs” at high redshifts, see more details in Appendix A). In contrast, the baryon SDGF $\mathcal{D}_{b}\left(k,z\right)$ is much smaller than $D\left(z\right)$ at high redshifts, while at low redshifts they are nearly the same due to the increasing coupling due to gravitation between the baryon and CDM fluids. This is demonstrated in Fig. 3, where BAO features are clearly manifest. Since $\mathcal{D}_{b}\left(k,z\right)\ll\mathcal{D}_{c}\left(k,z\right)$ at the dark ages, we have $\delta_{b}\left(\mathbf{x},z\right)\ll\delta_{c}\left(\mathbf{x},z\right)$ from Eq. (9), as illustrated in Fig 4. This has some important implications on the 21-cm signal at that epoch.
@@
Figure 4: Left (right) panel: realization of the CDM (baryon) over-density field at $z=300$.
II.2 Brightness temperature at the dark ages
[…]
The quantity $T_{s}$ that appears in Eqs. (11)-(12) is the spin temperature. This quantity is defined as $n_{1}/n_{0}\equiv 3\,\mathrm{exp}\left(-T_{\star}/T_{s}\right)$, where $n_{1}$ ($n_{0}$) denotes the number density of hydrogen atoms in the hyperfine triplet (singlet) state and $T_{\star}\equiv\hbar c/\left(k_{B}\lambda_{21}\right)\approx 68.2\,\mathrm{mK}$. In theory, the solution for $T_{s}$ has to be obtained by solving the Boltzmann equation for $n_{1}$ and $n_{0}$. Yet, we show in Appendix B that the solution for this equation is very well approximated by the well known expression for $T_{s}$ in equilibrium, even in the dark ages where $H\gg A_{10}$. This equilibrium expression reads
[…]
where $x_{e}\equiv n_{e}/\left(n_{\mathrm{H}}+n_{\mathrm{He}}\right)$ is the free-electron fraction, with $n_{e}$, $n_{\mathrm{H}}$ and $n_{\mathrm{He}}$ the free electron, hydrogen nuclei and helium nuclei number densities, respectively. The first line in Eq. (14) represents the change in temperature in the absence of external heating sources; the first term accounts for adiabatic cooling due to the expansion of the Universe, the second term accounts for adiabatic heating due to clustering of matter and the third term accounts for adiabatic cooling/heating as a result of increasing/lowering the total number of baryons in reionization/recombination. The second line in Eq. (14) represents heating by external sources, with the first (second) term accounting for heating by CMB (X-ray) photons. The expression for Compton heating by CMB photons is well-known and can be found in many papers in the literature, including Refs. Peebles (2020); Ma and Bertschinger (1995); Seager et al. (1999, 2000); Barkana and Loeb (2005b); Naoz and Barkana (2005). In contrast, the calculation of X-ray heating, relevant only during the cosmic dawn and reionization epochs, is much more complicated and model dependent, see e.g. Ref. Mesinger et al. (2011).
During the dark ages, the only relevant matter species that contributes to the 21-cm signal are the baryons, as they source the emission and absorption of 21-cm radiation by hydrogen atoms, the recombination rate that governs $T_{k}$, and the coupling of $T_{s}$ to $T_{k}$ via collisional excitations. Thus the fluctuations in the 21-cm signal during the dark ages are entirely proportional to $\delta_{b}$. Because $\delta_{b}\ll 1$ during the dark ages, the scale-dependent evolution of the baryon density field barely affects the global 21-cm signal, i.e. the sky-averaged brightness temperature as a function of redshift. This is demonstrated in Fig. 5, where the baryon over-density field is evolved with the SIGF (blue curve), or the SDGF (orange curve).
Figure 5: The 21-cm global signal as a function of redshift, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). The curves almost completely overlap each other.
where $\bar{T}_{21}$ is the global brightness temperature and $P_{21}\left(k,z\right)$ is the angle-averaged Fourier transform of the two-point correlation function $\langle\delta_{21}\left(\mathbf{x},z\right)\delta_{21}\left(\mathbf{x}^{\prime},z\right)\rangle$, while $\delta_{21}$ is the local contrast in the brightness temperature, $\delta_{21}\left(\mathbf{x},z\right)\equiv T_{21}\left(\mathbf{x},z\right)/\bar{T}_{21}\left(z\right)-1$. In a previous work Flitter and Kovetz (2024b), we have shown that during the dark ages $\Delta_{21}^{2}\left(k,z\right)\propto\mathcal{T}_{b}^{2}\left(k,z\right)\ll\mathcal{T}_{c}^{2}\left(k,z\right)$. The analysis in that work was based on the same SDGF formalism we present here, which we have shown to agree well with fully relativistic Boltzmann-Einstein solvers. In Fig. 6, we show $\Delta_{21}^{2}\left(k,z\right)$, as calculated by 21cmFirstCLASS after the incorporation of the SDGF in the code (we verify the correctness of our results in Appendix C). As expected, $\Delta_{21}^{2}\left(k,z\right)$ is suppressed during the dark ages when $\delta_{b}$ is correctly evolved with the SDGF.
Figure 6: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). For a zoomed-in version of the figure at cosmic dawn, see Fig. 15 in Appendix D.
III Cosmic dawn and reionization
III.1 Baryons
The role of the baryons in shaping the 21-cm signal is enhanced when the dark ages are over. As dark matter halos collapse, the gravitational pull that the baryons experience is increased. If the cooling timescale is shorter than the free-fall timescale, zero-metalicity popIII stars can form via gas accretion, mostly in molecular cooling galaxies (MCGs) Tegmark et al. (1997); Abel et al. (2002); Bromm and Larson (2004); Haiman and Bryan (2006); Trenti (2010); Plombat and Puy (2024). Baryonic feedbacks, such as supernovae Springel and Hernquist (2003); Wise and Abel (2008); Pawlik and Schaye (2009), photo-heating Hui and Gnedin (1997); Okamoto et al. (2008); Ocvirk et al. (2020); Katz et al. (2020) and Lyman-Werner radiation Bromm and Larson (2004); Fialkov et al. (2013); Haiman et al. (1997) hamper the star formation rate (SFR). Still, stars continue to form in more massive halos that host atomic cooling galaxies (ACGs) Qin et al. (2020, 2021a); Krumholz (2012). Those halos typically are assumed to contain mainly popII stars with low metalicity.
[…]
The local SFRD does not depend only on the local baryon density, but also on local velocities, specifically on the relative velocity between baryons and CDM, $\mathbf{v}_{cb}=\mathbf{v}_{c}-\mathbf{v}_{b}$. From Eqs. (2) and (4), we see that $\dot{\mathbf{v}}_{cb}=-H\mathbf{v}_{cb}$, which has the temporal solution $\mathbf{v}_{cb}\left(\mathbf{x},z\right)\propto\left(1+z\right)$. As for the spatial solution, adiabatic initial conditions imply Gaussian distributions for each one of the three components of $\mathbf{v}_{cb}$, and a Maxwell-Boltzmann distribution for $v_{cb}\equiv\left(\mathbf{v}_{cb}\cdot\mathbf{v}_{cb}\right)^{1/2}$. The variance of this distribution can be determined by defining $\theta_{cb}\equiv a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{cb}$, and then
[…]
Figure 7: The effect of the baryon-CDM relative velocity $v_{cb}$ on the 21-cm global signal. In the green curve, the effect of $v_{\mathrm{cb}}$ enters only through the modification of the minimum halo mass that can host popIII stars, as implemented in 21cmFirstCLASS. In the orange curve, small-scale suppression effects in $\sigma\left(M_{h},z\right)$ are also included, as in 21cmFAST, see more details in Sec. IV. The small extra delay between the orange and green curves can be seen more clearly in Fig. 1 (as the difference between the solid and dashed green curves at $z\gtrsim$10).
@@ III.1 Baryons
Once galaxies have started to form inside halos, the baryon density field is far from homogeneous as $\delta_{b}>1$ in regions of galaxy clusters. Thus, during cosmic dawn and reionization epochs linear perturbation theory is no longer applicable in determining $\delta_{b}\left(\mathbf{x},z\right)$. Non-linear perturbation theories, such as standard perturbation theory (or effective field theory) Baumann et al. (2012); Carrasco et al. (2012, 2014); Hertzberg (2014); Porto et al. (2014); Senatore (2015); Senatore and Zaldarriaga (2014); Baldauf et al. (2015), renormalized perturbation theory Crocce and Scoccimarro (2006); Bernardeau et al. (2008); Montesano et al. (2010), the renormalization group approach McDonald (2007), closure theory Taruya and Hiramatsu (2008), TimeRG theory Matarrese and Pietroni (2007); Pietroni (2008), or time-sliced perturbation theory Blas et al. (2016), must be used in order to find the “non-linear” perturbation in the baryon density field $\delta_{b}^{\mathrm{NL}}$.
In this work, we consider a 2nd-order Lagrangian perturbation theory Friedman and Schutz (1978); Buchert (1989); Moutarde et al. (1991); Buchert (1992, 1993); Bouchet et al. (1995); Hivon et al. (1995); Buchert (1995); Tatekawa (2005); Matsubara (2008); Carlson et al. (2009); Zheligovsky and Frisch (2014) (2LPT), as in 21cmFAST, though we improved its calculations in 21cmFirstCLASS to account for the scale-dependent growth of the baryons. Readers who wish to learn more about the scale-dependent 2LPT can find more details in Appendix D. The importance of evolving the baryon density field non-linearly is evident in Fig. 8; while linear perturbation theory seems to give a rough estimate for the distribution of baryonic matter on very large scales, it fails to capture the clustering of baryons at low redshifts and small scales, and it wrongly predicts a more homogeneous density field. When computing the 21-cm power spectrum, the imprecise calculation from linear perturbation theory yields an order unity error at $z\lesssim 8$, as demonstrated in Fig. 9. Importantly, this is the regime where upcoming 21-cm interferometers are mostly sensitive (c.f. Fig. 2).
Figure 8: Left (right) panel: realization of the baryon over-density field at $z=6$ when evolved linearly (non-linearly).
Figure 9: The 21-cm power spectrum as a function of redshift at $k=0.3\,\mathrm{Mpc}^{-1}$, when the baryon and the total matter over-density fields are evolved either linearly or non-linearly. The colors of the curves here match the same colors in Figs. 1-2, namely the blue, orange and green curves can be associated with 21cmFAST, Zeus21-like, and 21cmFirstCLASS, respectively. Here though, unlike Figs. 1-2, in all curves we distinguish between baryons and CDM (i.e. in all curves we use the SDGF), and we have the same $v_{cb}$ delaying effect (see differences in Fig. 7).
III.2 Total matter
During the cosmic dawn and reionization epochs, the 21-cm signal does not depend only on the baryon over-density $\delta_{b}$, but also on the CDM over-density field $\delta_{c}$, or more precisely on the total matter over-density $\delta_{m}$, Eq. (6). This is because the SFRD is given by the following expression Madau et al. (1997); Schneider et al. (2021); Muñoz et al. (2022); Muñoz (2023); Cruz et al. (2024b),
[…]
where $\dot{M}_{*}\left(\mathbf{x},M_{h},z\right)$ is the local stellar mass production rate in a halo of mass $M_{h}$ and $dn\left(\mathbf{x},M_{h},z\right)/dM_{h}$ is the conditional halo mass function (HMF); it is conditional in the sense that it depends on the value of the local total matter over-density $\delta_{m}\left(\mathbf{x},z\right)$. The mean of the conditional HMF is the “standard” HMF, defined as Lukic et al. (2007)
@@ III.2 Total matter
where $\mathcal{T}_{m}\left(k,z\right)\equiv\delta_{m}^{\mathrm{lin}}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the linear matter density transfer function and $W\left(kR\right)=3\left(\sin kR-kR\cos kR\right)/\left(kR\right)^{3}$. The function $f\left(\sigma\right)$ is the HMF fitting function and is calibrated on data from N-body simulations. Perhaps the simplest choice is to rely on the purely Gaussian fitting function from the Press-Schechter (PS) Press and Schechter (1974) HMF. In this work, we use instead the Sheth-Tormen (ST) HMF Sheth and Tormen (1999); Sheth et al. (2001); Sheth and Tormen (2002) (but see other popular mass functions in Refs. Jenkins et al. (2001); Warren et al. (2006); Reed et al. (2007); Tinker et al. (2008); Watson et al. (2013)).
In evaluating the integral in Eq. (17), the following quantity is often considered,
@@
This is the local collapsed fraction of halos more massive than $M_{\mathrm{min}}\left(z\right)$. The collapsed fraction is a very useful quantity since the extended Press-Schechter (EPS) theory predicts its value analytically Bond et al. (1991). According to EPS, for a PS HMF, in a large-scale region of comoving radius $R$, centered at $\mathbf{x}$, with a smoothed over-density $\delta_{m,R}$, the collapsed fraction of halos above $M_{\mathrm{min}}$ at redshift $z$ is
@@
Note that in writing Eq. (III.2) we have used the redshift-dependent $\sigma\left(M_{h},z\right)$ as given by Eq. (19), without imposing the SIGF approximation, $\sigma\left(M_{h},z\right)\propto D\left(z\right)$, which is less accurate at high redshifts and low masses (see more details in Appendix A).
IV Discussion
@@
One delicate nuance regarding the formalism presented above concerns the assumed matter density field that is used in Eq. (III.2). Press-Schechter theory Sheth and Tormen (1999), and the extended Press-Schechter theory Bond et al. (1991), are both established on linear cosmological perturbation theory. Hence, from a purely theoretical point of view, it would be more consistent to evaluate the smoothed $\delta_{m}$ that appears in Eq. (III.2), as well as $\sigma\left(M_{h},z\right)$, from linear theory. This is for example the underlying assumption in Zeus21, where the “linear” $\delta_{m}^{\mathrm{lin}}$ still results in non-linear fluctuations in the SFRD due to the erfc function in Eq. (III.2), an effect that can be captured by an exponential fit Muñoz (2023). In contrast, 21cmFAST uses the non-linear $\delta_{m}^{\mathrm{NL}}$ (from 2LPT calculations) in Eq. (III.2). The justification for this choice comes from better agreement with radiative transfer simulations at the epoch of reionization Zahn et al. (2011). [^4] We comment that the non-linear density field contains non-Gaussian statistics and therefore non-vanishing bispectrum, leading to a much more complicated $f_{\mathrm{coll}}$, see Eq. (5) in Ref. Lidz et al. (2013). Evidently, because the fluctuations in the brightness temperature are strongly connected to the fluctuations in the SFRD, the choice of working with either $\delta_{m}^{\mathrm{lin}}$ or $\delta_{m}^{\mathrm{NL}}$ has large implications on the resulting 21-cm power spectrum, as can be seen in Fig. 9. Our readers are hence forewarned.
The second nuance that we wish to highlight is the effect that $v_{cb}$ has on the local SFRD. As previously discussed, the local SFRD is suppressed in regions with large relative velocity. In 21cmFAST and 21cmFirstCLASS this effect is modelled in the code by increasing $M_{\mathrm{min}}$ in MCGs Muñoz et al. (2022),

### v2

@@ II.2 Growth factor
During the dark ages, the fluctuations in the matter density field were still much smaller than unity and linear perturbation theory can be applied in that epoch. After recombination took place, the linearized continuity and Euler equations for the baryon and CDM fluids on sub-horizon scales are Baumann (2022); Dodelson (2003),
@@
$\dot{\mathbf{v}}_{b}+H\mathbf{v}_{b}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ $-a^{-1}c_{s}^{2}\boldsymbol{\nabla}\delta_{b}$ (6)
@@
where $\delta_{c}$ is the CDM energy over-density, $\mathbf{v}_{c}$ is the proper CDM peculiar velocity, and $c_{s}$ the baryonic sound speed. Overdots denote derivatives with respect to the cosmological time $t$. Eqs. (5)-(8) are accompanied with the Poisson equation for the gravitational potential $\Phi$,
@@
Below the comoving Jeans scale, $\sim 10\,\mathrm{kpc}$, the fluctuations of the 21-cm signal are suppressed due to smoothing caused by the pressure of the baryonic gas Loeb and Zaldarriaga (2004); Naoz and Barkana (2005), as is manifest through the $c_{s}$ term in Eq. (6). We have showed in a previous work Flitter and Kovetz (2024b) that above the Jeans scale, at a wavenumber $k\lesssim 50\,\mathrm{Mpc}^{-1}$, the fluctuations in the 21-cm signal can be very well captured by omitting this term. Under this approximation, the equations of motion for the baryon and CDM fluids become identical and we can combine Eqs. (5)-(8) into two coupled second-order differential equations of the same form ($i=b,c$ for baryons and CDM),
@@
The CDM SDGF is nearly scale-invariant and can be very well approximated by $\mathcal{D}_{c}\left(k,z\right)\approx D\left(z\right)$, apart for BAO features that appear at $z\lesssim 1$ (while adopting the SIGF for the CDM density field causes “fake BAOs” at high redshifts, see more details in Appendix B). In contrast, the baryon SDGF $\mathcal{D}_{b}\left(k,z\right)$ is much smaller than $D\left(z\right)$ at high redshifts, while at low redshifts they are nearly the same due to the increasing coupling due to gravitation between the baryon and CDM fluids. This is demonstrated in Fig. 1, where BAO features are clearly manifest. Since $\mathcal{D}_{b}\left(k,z\right)\ll\mathcal{D}_{c}\left(k,z\right)$ at the dark ages, we have $\delta_{b}\left(\mathbf{x},z\right)\ll\delta_{c}\left(\mathbf{x},z\right)$ from Eq. (13), as illustrated in Fig 2. This has some important implications on the 21-cm signal at that epoch, as we show in Sec IV.1.
@@
Figure 2: Left (right) panel: realization of the CDM (baryon) over-density field at $z=300$. The color bar is the same for the two plots, hence highlighting the stronger fluctuations in the CDM field.
II.3 Non linear evolution
@@
Once galaxies have started to form inside halos, the baryon density field is far from homogeneous as $\delta_{b}>1$ in regions of galaxy clusters. Thus, during cosmic dawn and reionization epochs linear perturbation theory is no longer applicable in determining $\delta_{b}\left(\mathbf{x},z\right)$. Non-linear perturbation theories, such as standard perturbation theory (or effective field theory) Baumann et al. (2012); Carrasco et al. (2012, 2014); Hertzberg (2014); Porto et al. (2014); Senatore (2015); Senatore and Zaldarriaga (2014); Baldauf et al. (2015), renormalized perturbation theory Crocce and Scoccimarro (2006); Bernardeau et al. (2008); Montesano et al. (2010), the renormalization group approach McDonald (2007), closure theory Taruya and Hiramatsu (2008), TimeRG theory Matarrese and Pietroni (2007); Pietroni (2008), or time-sliced perturbation theory Blas et al. (2016), must be used in order to find the “non-linear” perturbation in the baryon density field $\delta_{b}^{\mathrm{NL}}$. In this work, we consider a 2nd-order Lagrangian perturbation theory Friedman and Schutz (1978); Buchert (1989); Moutarde et al. (1991); Buchert (1992, 1993); Bouchet et al. (1995); Hivon et al. (1995); Buchert (1995); Tatekawa (2005); Matsubara (2008); Carlson et al. (2009); Zheligovsky and Frisch (2014) (2LPT).
We provide in Appendix C the full derivation of the scale-dependent 2LPT equations. In brief, in this method parcels move from their initial Lagrangian coordinates $\mathbf{q}$ at redshift $z_{0}$ to their Eulerian coordinates $\mathbf{x}$ according the displacement vector field $\mathbf{s}$, such that $\mathbf{x}\left(\mathbf{q},z\right)=\mathbf{q}\left(z_{0}\right)+\mathbf{s}\left(\mathbf{q},z\right)$. The solution to $\mathbf{s}$ can then be expanded in perturbation theory, $\mathbf{s}=\mathbf{s}^{(1)}+\mathbf{s}^{(2)}$. Under the SIGF approximation, in which $\delta_{b}=\delta_{c}$, $\mathbf{s}^{(1)}$ and $\mathbf{s}^{(2)}$ are given by
$\mathbf{s}^{\left(1\right)}\left(\mathbf{q},z\right)$ $=$ $\left[D\left(z\right)-D\left(z_{0}\right)\right]\boldsymbol{\nabla}_{\mathbf{q}}\psi^{\left(1\right)}\left(\mathbf{q}\right)$ (15)
$\mathbf{s}^{\left(2\right)}\left(\mathbf{q},z\right)$ $=$ $\left[E\left(z\right)-E\left(z_{0}\right)\right]\boldsymbol{\nabla}_{\mathbf{q}}\psi^{\left(2\right)}\left(\mathbf{q}\right),$ (16)
where
$E\left(z\right)=-\frac{3}{7}D^{2}\left(z\right),\qquad\nabla_{\mathbf{q}}^{2}\psi^{(1)}\left(\mathbf{q}\right)=-\delta_{0}\left(\mathbf{q}\right),$ (17)
In Sec. III, we discuss how these equations change when the SDGF is introduced.
II.4 Conditional mass function
During the cosmic dawn and reionization epochs, the 21-cm signal does not depend only on the baryon over-density $\delta_{b}$, but also on the CDM over-density field $\delta_{c}$, or more precisely on the total matter over-density $\delta_{m}$, Eq. (10). This is because galaxies and stars, which emit crucial radiation that alters the signal, form inside collapsed dark matter halos. In analytical and semi-numerical models, the effect of the halos on the signal comes into play through the conditional halo mass function (HMF), $dn\left(\mathbf{x},M_{h},z\right)/dM_{h}$; it is conditional in the sense that it depends on the value of the local total matter over-density $\delta_{m}\left(\mathbf{x},z\right)$. The mean of the conditional HMF is the “standard” HMF, defined as Lukic et al. (2007)
@@ II.4 Conditional mass function
where $\mathcal{T}_{m}\left(k,z\right)\equiv\delta_{m}^{\mathrm{lin}}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the linear matter density transfer function and $W\left(kR\right)=3\left(\sin kR-kR\cos kR\right)/\left(kR\right)^{3}$. The function $f\left(\sigma\right)$ is the HMF fitting function and is calibrated on data from N-body simulations. Perhaps the simplest choice is to rely on the purely Gaussian fitting function from the Press-Schechter (PS) Press and Schechter (1974) HMF. While the simple form of the PS fitting function captures well the qualitative behavior of the HMF, it under-estimates the number of massive halos. For that reason, we use in this work instead the Sheth-Tormen (ST) HMF Sheth and Tormen (1999); Sheth et al. (2001); Sheth and Tormen (2002) (but see other popular mass functions in Refs. Jenkins et al. (2001); Warren et al. (2006); Reed et al. (2007); Tinker et al. (2008); Watson et al. (2013)).
The calculation of the emissivity of the radiating sources often involves the conditional halo mass function through the integral in Eq. (25). In evaluating that integral, the following quantity can be useful,
@@
This is the local collapsed fraction of halos more massive than $M_{\mathrm{min}}\left(z\right)$. From Eq. (21) it is clear then that the conditional HMF can be extracted from the collapsed fraction through the relation
$\frac{dn\left(\mathbf{x},z,M_{h}\right)}{dM_{h}}=-\frac{\bar{\rho}_{m,0}}{M_{h}}\frac{df_{\mathrm{coll}}\left(\mathbf{x},z;M_{h}\right)}{dM_{h}}.$ (22)
The collapsed fraction is a very useful quantity since the extended Press-Schechter (EPS) theory predicts its value analytically Bond et al. (1991). According to EPS, for a PS HMF, in a large-scale region of comoving radius $R$, centered at $\mathbf{x}$, with a smoothed over-density $\delta_{m,R}$, the collapsed fraction of halos above $M_{\mathrm{min}}$ at redshift $z$ is
@@
Note that in writing Eq. (II.4) we have used the redshift-dependent $\sigma\left(M_{h},z\right)$ as given by Eq. (20), without imposing the SIGF approximation, $\sigma\left(M_{h},z\right)\propto D\left(z\right)$, which is less accurate at high redshifts and low masses (see more details in Appendix B).
@@
One delicate nuance regarding the formalism presented above concerns the assumed matter density field that is used in Eq. (II.4). Press-Schechter theory Sheth and Tormen (1999), and the extended Press-Schechter theory Bond et al. (1991), are both established on linear cosmological perturbation theory. Hence, from a purely theoretical point of view, it would be more consistent to evaluate the smoothed $\delta_{m}$ that appears in Eq. (II.4), as well as $\sigma\left(M_{h},z\right)$, from linear theory. This is for example the underlying assumption in Zeus21, where the “linear” $\delta_{m}^{\mathrm{lin}}$ still results in non-linear fluctuations in the SFRD due to the erfc function in Eq. (II.4), an effect that can be captured by an exponential fit Muñoz (2023). In contrast, 21cmFAST uses the non-linear $\delta_{m}^{\mathrm{NL}}$ (from 2LPT calculations) in Eq. (II.4). The justification for this choice comes from better agreement with radiative transfer simulations at the epoch of reionization Zahn et al. (2011). [^5] We comment that the non-linear density field contains non-Gaussian statistics and therefore non-vanishing bispectrum, leading to a much more complicated $f_{\mathrm{coll}}$, see Eq. (5) in Ref. Lidz et al. (2013).
II.5 Star formation rate density
The role of the baryons in shaping the 21-cm signal is enhanced when the dark ages are over. As dark matter halos collapse, the gravitational pull that the baryons experience is increased. If the cooling timescale is shorter than the free-fall timescale, zero-metalicity popIII stars can form via gas accretion, mostly in molecular cooling galaxies (MCGs, other works often refer to molecular cooling halos for more generality Tegmark et al. (1997); Abel et al. (2002); Bromm and Larson (2004); Haiman and Bryan (2006); Trenti (2010); Plombat and Puy (2024)). Baryonic feedbacks, such as supernovae Springel and Hernquist (2003); Wise and Abel (2008); Pawlik and Schaye (2009), photo-heating Hui and Gnedin (1997); Okamoto et al. (2008); Ocvirk et al. (2020); Katz et al. (2020) and Lyman-Werner radiation Bromm and Larson (2004); Fialkov et al. (2013); Haiman et al. (1997) hamper the star formation rate (SFR). Still, stars continue to form in more massive halos that host atomic cooling galaxies (ACGs) Qin et al. (2020, 2021a); Krumholz (2012). Those halos typically are assumed to contain mainly popII stars with low metallicity, though popIII stars can still form in atomic cooling halos if there is sterilizing radiation background that prevents star formation earlier in the molecular cooling stage (when the halo was smaller and younger).
[…]
where $\dot{M}_{*}\left(\mathbf{x},M_{h},z\right)$ is the local stellar mass production rate in a halo of mass $M_{h}$. Clearly, the local SFRD is smaller in regions with smaller $\dot{M}_{*}$, while the latter quantity must be considerably low at halos of mass $\sim M_{\mathrm{min}}$, where the halo mass is insufficient for efficient cooling. In practice, Eqs. (22) and (II.4) are used in computing $\dot{\rho}_{*}\left(\mathbf{x},z\right)$ under the assumption of PS HMF, while the unconditional ST HMF is then used to normalize the mean SFRD Barkana and Loeb (2004, 2008); Mesinger et al. (2011).
The local SFRD does not depend only on the local baryon density, but also on local velocities, specifically on the relative velocity between baryons and CDM, $\mathbf{v}_{cb}=\mathbf{v}_{c}-\mathbf{v}_{b}$. From Eqs. (6) and (8), we see that above the Jeans scale $\dot{\mathbf{v}}_{cb}=-H\mathbf{v}_{cb}$, which has the temporal solution $\mathbf{v}_{cb}\left(\mathbf{x},z\right)\propto\left(1+z\right)$. As for the spatial solution, adiabatic initial conditions imply Gaussian distributions for each one of the three components of $\mathbf{v}_{cb}$, and a Maxwell-Boltzmann distribution for $v_{cb}\equiv\left(\mathbf{v}_{cb}\cdot\mathbf{v}_{cb}\right)^{1/2}$. The variance of this distribution can be determined by defining $\theta_{cb}\equiv a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{cb}$, and then
where $\mathcal{T}_{\mathrm{\theta_{cb}}}\left(k,z\right)=\theta_{cb}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the transfer function for $\theta_{cb}$. Evaluating this integral at recombination ($z\approx 1100$) reveals that baryons move with a mean relative velocity of $\sim 30\,\mathrm{km/sec}$ with respect to CDM, corresponding to a Mach number of 5 Tseliakhovich and Hirata (2010); Ali-Haïmoud et al. (2014). This supersonic velocity makes it harder for the baryons to collapse and form popIII stars in small dark matter halos Tseliakhovich et al. (2011); Dalal et al. (2010); Fialkov et al. (2012); Greif et al. (2011); Stacy et al. (2011); O’Leary and McQuinn (2012); Schauer et al. (2019); Hirano et al. (2018); Barkana (2016); Muñoz (2019a), thereby lowering the SFRD and delaying cosmic dawn.
[…]